OpenStudy (anonymous):

Find an exact value of cos (pi/12)

OpenStudy (ajprincess):

cos(pi/6)=cos(2*(pi/12)) =2cos^2(pi/12)-1 cos(pi/6)=cos30 =(sqrt3)/2 Does that help?